# Unified Cross-Modal Attention-Mixer Based Structural-Functional Connectomics Fusion for Neuropsychiatric Disorder Diagnosis

### v1

@@ II METHODOLOGY
Let’s consider the provided dataset as $C=\left\{\left(D_{n}^{s},D_{n}^{f}\right),W_{n}\right\}_{n=1}^{P}$ with a total of $P$ subjects where for each subject $n$, $D^{s}$ and $D^{f}$ denotes the structural and functional connectomics modality, respectively, while $W$ presents the associated label (00: HC or $1$: SZ). As illustrated in Fig. 1, our objective is to obtain a low-dimensional fused embedding for each $n_{th}$ subject across both connectomics modalities to perform improved graph-level classification task. Note: Throughout the following sections, the superscript denotes either modality to simplify notation.
@@ II-B2 Cross-modal Attention-Mixer Fusion
For leveraging both global and local interactions within the six cross-modal attention layers outputs, we employed the MLP-Mixer architecture in a multi-view setting to capture higher-level dependencies and population-wise disorder patterns that may not be fully addressed by cross-attention module alone. Given that we had three different input modalities: structural ($R^{s^{\prime}}$), functional ($R^{f^{\prime}}$), and the unified representation ($R^{u}$); we structured the outputs of these cross-modal attention layers based on three distinct views. As illustrated in Fig. 1, the outputs of the cross-attention layers, where the $Q$ and $K$ matrices were derived from the same modality, were concatenated and fed into a seperate MLP-Mixer layer. Subsequently, the outputs of the multi-view MLP-Mixer layers: $H_{1},H_{2},H_{3}$ were concatenated and passed through another MLP-Mixer layer to obtain the final representation $H_{c}$
@@
SC N - - 71.43$\pm$1.46 71.33$\pm$2.04 66.00$\pm$1.81 74.29$\pm$1.74 78.72$\pm$2.22 74.33$\pm$1.12
SC Y - - 76.00$\pm$1.79 74.17$\pm$2.35 70.40$\pm$1.90 78.85$\pm$1.26 80.20$\pm$2.10 78.47$\pm$1.38
FNC N - - 73.33$\pm$1.94 75.17$\pm$2.35 71.93$\pm$2.06 76.19$\pm$1.36 79.54$\pm$1.72 75.24$\pm$1.51
FNC Y - - 77.14$\pm$1.67 77.33$\pm$2.13 75.87$\pm$1.42 80.67$\pm$1.15 81.67$\pm$2.47 77.57$\pm$1.42
SC+FNC Y N Concat 78.38$\pm$1.80 77.83$\pm$1.90 77.46$\pm$1.66 80.75$\pm$1.83 81.90$\pm$2.33 77.71$\pm$1.10
SC+FNC Y N DCCA 80.26$\pm$1.65 78.48$\pm$2.34 79.13$\pm$1.75 81.33$\pm$1.21 82.71$\pm$1.31 80.68$\pm$1.29
SC+FNC Y N Cross-Att 84.55$\pm$1.26 82.43$\pm$1.65 80.48$\pm$1.71 82.16$\pm$1.18 84.71$\pm$1.74 80.45$\pm$1.48
SC+FNC Y N ConneX 84.67$\pm$1.67 84.83$\pm$1.94 81.35$\pm$1.54 82.84$\pm$1.81 84.77$\pm$1.76 81.86$\pm$1.91
SC+FNC Y Y Concat 80.55$\pm$1.58 80.36$\pm$1.22 79.68$\pm$1.11 80.88$\pm$0.97 83.18$\pm$2.15 81.39$\pm$1.16
SC+FNC Y Y DCCA 83.92$\pm$1.13 84.58$\pm$2.05 82.37$\pm$1.03 81.67$\pm$2.49 83.67$\pm$2.42 81.33$\pm$2.16
SC+FNC Y Y Cross-Att 85.34$\pm$1.18 86.39$\pm$1.11 84.57$\pm$1.13 83.01$\pm$1.35 84.91$\pm$1.74 82.41$\pm$1.23
SC+FNC Y Y ConneX 88.53$\pm$1.49 93.33$\pm$1.33 85.56$\pm$1.29 85.71$\pm$1.11 86.67$\pm$1.05 84.05$\pm$1.25

### v2

@@ II METHODOLOGY
Let’s consider the provided dataset as $C=\left\{\left(D_{n}^{s},D_{n}^{f}\right),W_{n}\right\}_{n=1}^{P}$ with a total of $P$ subjects where for each subject $n$, $D^{s}$ and $D^{f}$ denotes the structural and functional connectomics modality, respectively, while $W$ presents the associated label (0: HC or $1$: SZ). As illustrated in Fig. 1, our objective is to obtain a low-dimensional fused embedding for each $n_{th}$ subject across both connectomics modalities to perform improved graph-level classification task. Note: Throughout the following sections, the superscript denotes either modality to simplify notation.
@@ II-B2 Cross-modal Attention-Mixer Fusion
For leveraging both global and local interactions within the six cross-modal attention layers outputs, we employed the MLP-Mixer architecture in a multi-view setting to capture higher-level dependencies and population-wise disorder patterns that may not be fully addressed by cross-attention module alone. Given that we had three different input modalities: structural ($R^{s^{\prime}}$), functional ($R^{f^{\prime}}$), and the unified representation ($R^{u}$); we structured the outputs of these cross-modal attention layers based on three distinct views. As illustrated in Fig. 1, the outputs of the cross-attention layers, where the $Q$ and $K$ matrices were derived from the same modality, were concatenated and fed into a seperate MLP-Mixer layer. Subsequently, the outputs of the multi-view MLP-Mixer layers: $H_{1},H_{2},H_{3}$ were concatenated and passed through another MLP-Mixer layer to obtain the final representation $H_{c}$ .
@@
[…]